# Opinion Clustering under the Friedkin-Johnsen Model: Agreement in Disagreement

### v2

@@ 4 Topology-based opinion clustering
An agent $p\in\mathcal{V}$ is called an LTP agent if the following hold:
@@
    additionally, if $p\in\mathcal{V}\setminus\mathcal{I}$, then there does not exist an agent $c\in\mathcal{V}$ such that all paths from each agent $s\in\mathcal{I}_{p}$ to $p$ traverse $c$.
The non-influential agent $q$ is then said to be persuaded by $p$.
The set of all non-influential agents persuaded by an LTP agent $p$ is denoted by $\mathcal{N}_{p}$. If $p\in\mathcal{I}$ is an influential agent, condition (i) is sufficient to ensure that $p$ is an LTP agent. The condition (ii) in Defn. 2 is applicable only if $p$ is a non-influential agent. It conveys that a non-influential agent $p$, persuaded by an LTP agent $c$ (i.e $p\in\mathcal{N}_{c}$), is not an LTP agent itself. The following example illustrates which agents are LTP agents and which are the agents it persuades.
@@
Consider the network $\mathcal{G}$ shown in Fig. 1(a). Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential. Fig 2(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses $3$. Further, the paths from $6$ to $4$ (and $5$) also traverse $3$ as shown in Fig. 2(b). Since all paths from $2$ and $6$ to $3$ do not traverse any common node. Thus, $3$ is an LTP agent and $\mathcal{N}_{3}=\{4,5\}$. Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
Note that each path from $2$ and $6$ to $5$ traverses $4$. Hence, $4$ satisfies condition 1) in Definition 2 However, since $4\in\mathcal{N}_{3}$, it is not an LTP agent.
@@
Figure 2: Each path from the stubborn agents to 44 and 55 traverses $3$.
Remark 2
The notion of an LTP agent is analogous to that of cut vertices in undirected graphs (and strong articulation points in digraphs). A node $v$ is a cut vertex (or a strong articulation point in a strongly connected digraph) if and only if there are distinct nodes $x$ and $y$ such that every path from $x$ to $y$ traverses $v$ [19]. Since LTP agents are defined for digraphs, each LTP agent in a strongly connected digraph is a strong articulation point; however, the converse does not hold (e.g., node $4$ in Example 2).
The previous discussion describes the properties of the LTP agents. In larger networks, however, multiple LTP agents may exist. This leads to the question: Can a non-influential agent be persuaded by more than one LTP agent?
@@
Under C3, there exists a path $P_{1}$ from $s$ to $q$ that traverses $i$ before $j$ and another path $P_{2}$ from $u$ to $q$ that traverses $j$ before $i$. We can represent $P_{1}$ as $s\to i\to j\to q$ and $P_{2}$ as $u\to j\to i\to q$. By definition, a path does not contain repeated nodes. It follows from $P_{1}$ that a path $j\to q$ exists that does not traverse $i$. Further, we know from $P_{2}$ that a path from $u\to j$ does not traverse $i$. Therefore, a path from $u\to q$ exists that does not traverse $i$. Consequently, $q\notin\mathcal{N}_{i}$ which contradicts our assumption that $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$. Since all of the conditions: C1, C2 and C3, lead to a contradiction, the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.
@@
Lemma 2 establishes that an LTP agent and the agents it persuades form a disjoint group in the network. Certain agents in $\mathcal{G}$ can exist who are neither an LTP agent nor are persuaded by any LTP agent. Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents.
4.1 The role of network topology on opinion formation
In this subsection, we present a framework that relates the final opinions of agents with the underlying network topology. Subsequently, we employ this framework to demonstrate that the topological relation between an LTP agent and the agent(s) it persuades leads to the formation of an opinion cluster.
Consider a network $\mathcal{G}$ of $n$ agents with opinions evolving under the FJ model (1). Let $\mathcal{G}$ have $m$ stubborn agents. Then, the final opinions satisfy: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$. We can rewrite these linear equations as follows:
$\underbrace{\begin{bmatrix}(I-(I-\beta)W)&-\eta&\\ 0&0&\end{bmatrix}}_{R}\ \underbrace{\begin{bmatrix}\mathbf{x}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}}_{z}=\mathbb{0}$ (2)
where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the $m$ stubborn agents. The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$. Henceforth, we represent eqn. (2) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
Remark 4.2.
The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$. The associated graph $\mathcal{G}(R)$ has $n+m$ nodes. Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents. Each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$. The source $n+i$ has only one outgoing edge to the node associated with the final opinion of the corresponding stubborn agent. For example, in Fig. 3, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has outgoing edge $(n+1,2)$.
@@ 4.1 The role of network topology on opinion formation
Consider the network $\mathcal{G}$ in Fig. 1(a) with stubborn agents $2$ and $6$. The corresponding network $\mathcal{G}(R)$, shown in Fig. 3, is derived from eqn. (2). Each node associated with the initial opinion of a stubborn agent is shown in turquoise.
Figure 3: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
Remark 4.3.
Under Kron Reduction, we determine the Schur complement (4) of the loopy Laplacian matrix $Q$. Since $R$ is a Laplacian matrix, it can be reduced by Kron Reduction for a suitable choice of the subset of nodes $\alpha$. If $\mathcal{G}$ does not have any oblivious agents, then each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$. Thus, by Statement 1) in Lemma 5.1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$.
On the contrary, if $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path in $G(R)$ to this agent. For such networks, the set $\alpha$ must include the oblivious agents belonging to each iSCC and sources $\{n+1,...,n+m\}$. For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$. Thus, Statement 1) in Lemma 5.1 holds.
Consider a set of linear equations $M\mathbf{y}=\mathbb{0}$, where $\mathbf{y}\in\mathbb{R}^{p}$. Eliminating $y[\alpha^{c}]$ using Schur complement (4), yields the reduced system of equations: $M/\alpha^{c}\ \mathbf{y}[\alpha]=0$ (Pg. 18 in [21]). Similarly, we eliminate the states $z[\alpha^{c}]$ in the linear eqns. (2) using the Schur Complement to obtain,
where $\omega=\alpha\setminus\{n+1,...,n+m\}$. Thus, the Kron Reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents.
4.2 Main Results
Using the Kron Reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
Theorem 1.
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents. Let Assumption 1 hold. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1). If $p$ is an LTP agent, then $p$ and the nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
The proof is given in the Appendix.
In [15], the authors show that the final opinion of an agent depends on the paths (the weights of edges on the path) to it from the stubborn agents. However, it remains unclear what topological conditions on the paths from stubborn agents to distinct agents ensure that the final opinions of these agents converge to the same value. Theorem 1 addresses this gap by proving that the topological relation between an LTP agent and the agents it persuades leads to the formation of an opinion cluster. A major advantage of this result is that regardless of edge weights in the network $\mathcal{G}$ and the stubbornness values, an LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster. Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ and agents in $\mathcal{N}_{3}=\{4,5\}$ form an opinion cluster. Similarly, the LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$.
Remark 4.4.
In [12], the authors derived the conditions to obtain $k\in\mathbb{N}$ opinion clusters in weakly connected digraphs, under the assumption that each cycle in the digraph has at most one stubborn agent. Theorem 1 generalises these conditions to form opinion clusters in any arbitrarily connected digraphs, thereby removing this restrictive assumption.
Moreover, in [12] an opinion cluster is formed by group of agents composed either (1) of a single stubborn agent with the rest being non-stubborn, or (2) of all the oblivious agents in the network $\mathcal{G}$ when $\mathcal{G}$ has a unique aperiodic iSCC of oblivious agents. Hence, a group containing only non-influential agents does not form any opinion cluster in [12]. In contrast, Theorem 1 establishes that an opinion cluster can be formed by any group of non-influential agents if one of them is an LTP agent $p$ and the rest are in $\mathcal{N}_{p}$ (e.g., the agents $3,4,5$ in Example 1).
Remark 4.5.
Consider an agent $p\in\mathcal{V}$. To verify if $p$ is an LTP agent, we first remove all the outgoing edges of $p$, which takes $O(n)$ time. Next, the modified $\mathcal{G}$ is traversed $m+t$ times using the Depth First Search (DFS), once from each of the $m$ stubborn agents and the $t$ iSCCs composed of oblivious agents. As one iteration of DFS has complexity $O(n+|\mathcal{E}|)$, this step requires $O((m+t)(n+|\mathcal{E}|))$ time, where $|\mathcal{E}|$ denotes the number of edges. If a node remains unvisited after all these traversals, then $p$ satisfies Condition (i) in Defn 2. Let $\mathcal{U}_{p}$ be the set of unvisited nodes. By repeating this step for all $p\in\mathcal{V}$, we identify all agents that satisfy Condition (i) and store them as $\mathcal{H}$. This step has a time complexity of $O((t+m)n(n+|\mathcal{E}|))$ which reduces to $O(n(n+|\mathcal{E}|))$ since $t+m\ll n$. If $\mathcal{H}\subseteq\mathcal{I}$, identifying all LTP agents requires $O(n(n+|\mathcal{E}|))$ time.
Next, we verify condition (ii) in Defn. 2 for each non-influential agent in $\mathcal{H}$. For an non-influential agent $p$, we check if there is a $c\in\mathcal{H}\setminus\{p\}$ such that $p\in\mathcal{U}_{c}$. This step requires $O(n^{2})$ time. Repeating it for all $p\in\mathcal{H}$ yields a worst-case complexity of $O(n^{3})$. Thus, the overall time complexity of identifying all LTP agents in $\mathcal{G}$ is $O(n^{3})$.
Using the topological conditions in Theorems 1, a network can be designed to obtain a predefined number of opinion clusters with each cluster formed by a desired set of agents.
Corollary 2.
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents and $t\geq 0$ iSCCs with oblivious agents such that Assumption 1 holds. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1). Let $\mathcal{O}$ be the set of the oblivious agents that either belong to an iSCC or have paths from only one iSCC in $\mathcal{G}$. Then, $k\geq t+m$ opinion clusters form in $\mathbf{x}^{*}$ if the agents in set $\mathcal{V}\setminus\mathcal{O}$ can be partitioned into ktk-t disjoint groups with each group composed of an LTP agent and the nodes persuaded by it.
The proof of Corollary 2 is given in the Appendix. The following example illustrates the conditions presented in Corollary 2 to obtain the desired opinion clustering.
Example 4
@@ 4.2 Main Results
Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig. 4(a). The agents $1,2$ and $3$ are oblivious agents which together form an iSCC (highlighted by a dashed red circle); the remaining agents are non-oblivious. The non-oblivious agents in $\mathcal{G}$ form $4$ disjoint groups highlighted by the green dashed boxes in Fig. 4(a). Each group has an LTP agent (the node with green boundary) and the remaining agents in the group are persuaded by it. For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form $5$ opinion clusters as shown in Fig. 4(b). One cluster is formed by the oblivious agent belonging to the iSCC and the rest are formed by the agents belonging to the same group.
@@
5 Conclusions
The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models. Due to varied biases, even the individuals in a closely connected group hold diverse opinions. Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster. This work examines the role of the underlying network topology in the emergence of these clusters. To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades. Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent. Using Kron Reduction, we establish in Theorem 1 that an LTP agent and the set of agents it persuades form an opinion cluster in any arbitrarily-connected digraph under the FJ framework. Some additional key insights from our results are as follows: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents. (ii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [12] to any arbitrary digraphs. (iii) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering.
In future, we plan to use the LTP agents for desirably shaping the final opinions in real-world social networks.
Appendix
Kron Reduction: Let $M\in\mathbb{R}^{p\times p}$ with index set $\alpha\subseteq[p]$ and $\alpha^{c}:=[p]\setminus\alpha$. If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
Under Kron Reduction, the Schur complement of the loopy Laplacian matrix $Q$ is evaluated.
Lemma 5.1 ([18]).
Consider a loopy Laplacian matrix $Q\in\mathbb{R}^{p}$ with $\alpha\subset[p]$. Let $\mathcal{G}(Q)$ be the digraph associated with $Q$.
  1. 1.
    $Q/\alpha^{c}$ is well defined ( i.e. $Q[\alpha^{c}]$ is invertible) if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$. 111(Note that we adapt this condition to our framework because we consider $w_{ij}>0$ when $(j,i)\in\mathcal{E}$.)
  2. 2.
5.1 Proof of Theorem-1
Before moving to the proof of Theorem-1, the following Lemma needs to be stated.
Lemma 5.2.
Consider a weakly connected network $\mathcal{G}$ with $m\geq 1$ stubborn agents and matrix $R$ derived using eqn. (2). Let $p$ be an LTP agent and $q\in\mathcal{N}_{p}$. If $\alpha$ is the union of set $\{p,q,n+1,....,n+m\}$ and all the oblivious agents in iSCC in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$. Then,

### v3

@@ V Topology-based opinion clustering
An agent $p\in\mathcal{V}$ is called an LTP agent if the following holds:
@@
    additionally, if $p\in\mathcal{V}\setminus\mathcal{I}$, an agent $c\in\mathcal{V}$ such that all paths from each agent $s\in\mathcal{I}_{p}$ to $p$ traverse $c$ must not exist.
For an LTP agent $p$, the non-influential agent $q$ satisfying condition (i) is said to be persuaded by $p$.
The set of all non-influential agents persuaded by an LTP agent $p$ is denoted as $\mathcal{N}_{p}$. Note that if $p\in\mathcal{I}$ is an influential agent, condition (i) alone ensures that $p$ is an LTP agent. The condition (ii) in Defn. 2 is required to be satisfied only when $p$ is non-influential. It implies that if a non-influential agent $p$ is persuaded by an LTP agent $c$ (i.e $p\in\mathcal{N}_{c}$), then $p$ itself cannot be an LTP agent.
The following example illustrates the identification of an LTP agent and the agent(s) it persuades.
@@
Consider the network $\mathcal{G}$ shown in Fig. 1(a). Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential and have a path to all agents. Fig 3(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses through $3$. Moreover, the paths from $6$ to $4$ (and $5$) also traverse through $3$ as displayed in Fig. 3(b). Since the paths from $2$ and $6$ to $3$ do not traverse any common node. Thus, $3$ is an LTP agent and $\mathcal{N}_{3}=\{4,5\}$. Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
Note that each path from $2$ and $6$ to $5$ traverses $4$. Hence, $4$ satisfies condition 1) in Defn. 2. However, since $4\in\mathcal{N}_{3}$, it is not an LTP agent.
@@
Figure 3: Node 33 is an LTP agent and it persuades 44 and $5$ as each path from the stubborn agents traverse $3$.
Suppose a network $\mathcal{G}$ has multiple LTP agents. The following question naturally arises: Can an agent be persuaded by two different LTP agents? The following Lemma presents an answer to this question.
@@
Under C3, there exists a path $P_{1}$ from $s$ to $q$ that traverses $i$ before $j$ and another path $P_{2}$ from $u$ to $q$ that traverses $j$ before $i$. We can represent $P_{1}$ as $s\to i\to j\to q$ and $P_{2}$ as $u\to j\to i\to q$. By definition, a path does not contain repeated nodes. It follows from $P_{1}$ that a path $j\to q$ exists that does not traverse $i$. Further, we know from $P_{2}$ that a path from $u\to j$ does not traverse $i$. Therefore, a path from $u\to q$ exists that does not traverse $i$. Consequently, $q\notin\mathcal{N}_{i}$ which contradicts our assumption that $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$. Since all of the conditions: C1, C2 and C3, lead to a contradiction, it means that the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold. ∎
@@
Lemma 3 establishes that an LTP agent and the agents it persuades form a disjoint group in the network. However, there can be certain agents who are neither an LTP agent nor are persuaded by any LTP agent. Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents. The following results presents the relation between the final opinions of agents within each such group.
Theorem 1
Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2). If each iSCC in $\mathcal{G}$ composed of the oblivious agents is aperiodic, then final opinions of an LTP agent $p$ and nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
The proof of Theorem 1 is given in the Appendix.
Theorem 1 establishes that an LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster in the final opinion. Note that the existence of an LTP agent in a network depends only on paths from the influential agents; hence, it is a purely graph topological property. Consequently, an LTP agent $p$ and $\mathcal{N}_{p}$ form an opinion cluster independent of the edge weights and the stubbornness of the stubborn agents.
Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ forms an opinion cluster with agents in $\mathcal{N}_{3}=\{4,5\}$ and LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$. Additionally, it explains the formation of a new opinion cluster for the modified graph $\mathcal{\hat{G}}$. This occurs because the addition of edge $(1,5)$ results in a path from $6$ to $5$ that does not traverse $3$. Thus, $5\notin\mathcal{N}_{3}$ and it does not form an opinion cluster with $3$ and $4$, as shown in Fig. 1(d).
By definition, if an LTP agent $p$ is oblivious, then the agents in $\mathcal{N}_{p}$ are also oblivious. Consequently, it follows from Theorem 1 that a set of non-influential oblivious agents can also form an opinion cluster. Since the opinion of an oblivious agent evolves according to DeGroot model and is not affected by stubborn behaviour, an LTP agent ensures the formation of opinion clusters even in DeGroot framework. Note that since nodes in $\mathcal{N}_{q}$ are always non-influential, the desired opinion clusters can be formed in DeGroot framework only in weakly connected networks.
Remark 3
In [15], a network with at most one stubborn agent in each cycle is shown to form opinion clusters under the FJ model. Theorem 1 generalises the existing results by presenting LTP-based conditions to form an opinion cluster in any arbitrary digraph.
Further, in [15], each opinion cluster is formed by a group that contains at least one influential agent: a stubborn agent with non-stubborn agents or the oblivious agents in an iSCC. In contrast, Theorem 1 demonstrates that a group composed only of non-influential agents can also form an opinion cluster provided that one of them is an LTP agent and the rest are persuaded by it.
Next, we illustrate the topological conditions presented in Theorem 1 using the following example.
@@
Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig. 4(a). The agents $1,2$ and $3$ are oblivious agents which together form an iSCC (highlighted by a dashed red circle); the remaining agents are non-oblivious. The non-oblivious agents in $\mathcal{G}$ form $4$ disjoint groups highlighted by the green dashed boxes in Fig. 4(a). Each group has an LTP agent (the node with green boundary) and the remaining agents in the group are persuaded by it. For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form $5$ opinion clusters as shown in Fig. 4(b). One cluster is formed by the oblivious agents in the iSCC and the rest are formed by the $4$ disjoint groups of agents; each composed of an LTP agent and those persuaded by it.
@@
VI Conclusions
The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models. Due to varied biases, even the individuals in a closely connected group hold diverse opinions. Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster. This work examines the role of the underlying network topology in the emergence of these clusters. To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades. Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent. Using Kron reduction, we establish in Theorem 1 that the final opinions of an LTP agent and the set of agents it persuades are equal. Thus, they form an opinion cluster in any arbitrarily-connected digraph under the FJ framework. Interestingly, such LTP agents lead to the formation of opinion clusters even under the DeGroot model (see Theorem 1).
Some additional key insights from our results are: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents. (ii) Notably, an LTP agent may or may not be influential, but still shapes the final opinions of the agents it persuades. (iii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [15] to any arbitrary digraphs. (iv) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering as demonstrated in Example 4.
In future, we plan to examine the potential of LTP agents in desirably shaping the final opinions in social networks.
VII Appendix
Before moving to the proof of Theorem 1, the following Lemma needs to be stated.
Lemma 4
Consider a weakly connected network $\mathcal{G}$ with $m\geq 1$ stubborn agents and matrix $R$ derived using eqn. (3). Let $\alpha$ be a set containing $\{n+1,....,n+m\}$ and all the oblivious agents in the iSCCs in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$. Then,